# Active robustness against the detuning-error for Rydberg quantum gates

### v1

@@ Abstract
Error suppression to the experimental imperfections is a central challenge for useful quantum computing. Recent studies have shown the advantages of using single-modulated pulses based on optimal control which can realize high-fidelity two-qubit gates in neutral-atom arrays. However, typical optimization only minimizes the ideal gate error in the absence of any decay, which allows the gate to be passively influenced by all error sources leading to an exponential increase of sensitivity when the error becomes larger. In the present work, we propose the realization of two-qubit $C_{Z}$ gates with active robustness against two-photon detuning errors. Our method depends on a modified cost function in numerical optimization for shaping gate pulses, which can minimize, not only the ideal gate error but also the fluctuations of gate infidelity over a wide error range. We introduce a family of Rydberg blockade gates with active robustness towards the impacts of versatile noise sources such as Doppler dephasing and ac Stark shifts. The resulting gates with robust pulses can significantly increase the insensitivity to any type of errors acting on the two-photon detuning, benefiting from a relaxed requirement of colder atomic temperatures or more stable lasers for current experimental technology.
@@ I Introduction
The ability to achieve high-fidelity two-qubit quantum gates is a key task for large-scale quantum computing with trapped neutral atoms [1, 2, 3, 4], while noises from versatile experimental imperfections will fundamentally limit the achievable gate fidelity [5], leading to a huge gap to other state-of-the-art platforms such as superconducting qubits, trapped ions, quantum dots and nuclear magnetic resonance systems [6, 7, 8, 9, 10]. Recent work has demonstrated that the execution of modulated pulses based on optimal control can realize high-fidelity entangling operations [11]. And to date the best reported two-qubit gate fidelity using optimal control and atomic dark state reaches $\mathcal{F}=0.995$ [12]. The optimal control technique enables a smooth modulation of the laser amplitude or phase in time without the requirement for individual addressability of atoms [13, 14, 15, 16, 17, 18]. This outperforms the original $\pi$-gap-$\pi$ method [19, 20, 21] by suppressing severe decoherence between the ground and Rydberg states [21, 22, 23].
@@
The typical optimization focuses on minimizing the ideal gate infidelity in the absence of any noise. e.g. A time-optimal gate contains the merit of minimizing the gate duration together with a smaller time-spent in the Rydberg state which makes the gate robust to intrinsic Rydberg decay error [24]. Whereas other technical errors will passively affect the gate performance suggesting the necessity of subsequent error correction methods regardless of the system [25, 26, 27, 28, 29, 30]. Recently, performing low-error quantum gates with noise-robust optimization has become an exciting frontier. To alleviate the sensitivity of gates to type of errors that occur, e.g. constructing a specific cost function that possess optimal robustness with respect to some errors, is more attractive [31, 32, 33]. Via utilizing quantum optimal control with GRAPE algorithm [34] a family of Rydberg blockade gates with significantly improved performance is presented that can be simultaneously robust against the amplitude inhomogeneity and Doppler shifts [35]. Although these works may open unique opportunities for noise-robust quantum computing with atomic qubits, preserving an active robustness of systems to type of errors for Rydberg gates is still a big challenge.
@@
Figure 1: Two-qubit $C_{Z}$ gate protocol. (a) Atom-level diagram. Qubits are encoded in two hyperfine clock states $|0\rangle$ and $|1\rangle$ of ${}^{87}\rm{Rb}$ atoms. Atomic excitation to the Rydberg state $|r\rangle$ is carried out by a two-photon transition via a largely-detuned intermediate state $|p\rangle$(detuned by $\Delta_{0}$), driven by 420 nm and 1013 nm lasers. $|d\rangle$ is a leakage state that collects most of the spontaneous decays from $|p\rangle$ and $|r\rangle$ except returning to $|0\rangle$ and $|1\rangle$. The branching radios used for different scattering rates are: $b_{r\to 0}=0.059$, $b_{r\to 1}=0.055$, $b_{r\to d}=0.886$, and $b_{p\to 0}=0.1354$, $b_{p\to 1}=0.2504$, $b_{p\to d}=0.6142$, together with a $3751/\gamma_{r}=$ 375 $\mu$s lifetime for $|r\rangle$ and a $0.1181/\gamma_{p}=$ 0.118 $\mu$s lifetime for $|p\rangle$, chosen from [36]. The van der Waals interaction leads to a large energy shift $B$ if both atoms are in $|r\rangle$ preventing their simultaneous excitation in the Rydberg blockade limit. $\epsilon_{\Delta}$ and $\epsilon_{\delta}$ represent unknown detuning errors respectively to the intermediate and Rydberg states possibly from versatile experimental imperfections. (b) The pulse sequence $\Omega_{p}(t),\Omega_{c}(t)$ and the non-fluctuated intermediate detuning $\Delta_{0}$. Note that $\Delta_{0}$ is switched between two STIRAP pulses. (c) The gate infidelity $1-{\mathcal{F}_{der(to)}}$ in the absence of spontaneous decay as a function of the two-photon detuning error $\epsilon_{\delta}$, given by the DER and TO pulses.
@@
In this work, we present the realization of two-qubit Rydberg $C_{Z}$ gates with active robustness to all errors associated with the variation in two-photon detuning [37, 38]. To implement the noise-robust optimization, we mainly modify the cost function in the numerical Genetic algorithm which minimizes not only the ideal gate error but also the infidelity fluctuation over a wide error range, allowing the optimal pulse to be actively robust against the variation of errors. In addition, we adopt a realistic model by considering a native two-photon transition using adiabatic double-STIRAP pulse [39] which has shown a better robustness to the variation of experimental parameters such as laser amplitude and pulse shape [40, 41, 42]. To characterize the resulting two-qubit $C_{Z}$ gates , we comparably exploit the results from typical and noise-robust optimizations. The latter can produce experimentally realizable pulses that are more robust to the variation of two-photon detunings caused by Doppler dephasing from finite atomic temperature [43] and ac Stark shifts [44, 45]. Our approach can be generalized to other type of errors paving the avenue for realizing noise-robust quantum gates in future experiments.
@@
Finally we admit that there are two main weakness of this protocol. First, the modified cost function contains two competing terms that means the improvement of robustness should be at the expense of a high ideal gate fidelity (see Appendix B). For a noise-robust pulse, we achieve the gate with an ideal fidelity of $\mathcal{F}=0.9971$. Second, the proposed gates will be affected by other error sources not associated to the two-photon detuning. As compared to the typical-optimal pulse, our pulses are not explicitly robust to those errors. Especially due to the use of an adiabatic pulse that elongates the gate duration, our scheme is more susceptible to the fluctuation in laser phase [46, 47].
II Model description and Hamiltonian
@@
Consider a level scheme for two-qubit $C_{Z}$ gates, as shown in Fig.1(a). Each atom is modeled as a native four-level system with long-lived hyperfine qubit states $|0\rangle,|1\rangle$, an intermediate state $|p\rangle$ and an uppermost Rydberg state $|r\rangle$ [48]. The coherent excitation from $|1\rangle$ to $|r\rangle$ is mediated by a resonant two-photon transition with laser Rabi frequencies $\Omega_{p},\Omega_{c}$, largely detuned by $\Delta_{0}$ from the intermediate state $|p\rangle$, i.e. $|\Delta_{0}|\gg\Omega_{p},\Omega_{c}$. We assume a strong Rydberg blockade condition where two atoms trapped in two optical tweezers [49] with a distance smaller than the blockade radius, can bring a strong energy shift $B$ to prevent simultaneous excitation of both atoms [51, 52].
@@
for single control or target qubit. Here, $\epsilon_{\Delta}$ and $\epsilon_{\delta}$ treats as the unknown yet constant detuning error over the gate duration possibly caused by versatile noise sources such as atomic velocity and laser fluctuations. We point out that, the intermediate detuning error $\epsilon_{\Delta}$ is less important because $|\Delta_{0}|\gg\epsilon_{\Delta}$ arising a negligible impact $10^{-9}\sim 10^{-6}$ on fidelity(see Appendix C) [53]. Nevertheless, the laser frequency is resonant with the two-photon transition between $|1\rangle$ and $|r\rangle$, making the role of two-photon detuning error $\epsilon_{\delta}$ more crucial in the scheme.
@@ III.1 Double STIRAP pulse sequences
We proceed by describing the implementation of specific pulse sequences that result in a $\pi$-phase($C_{Z}$) gate. The requirement for a universal controlled-phase gate is that the laser pulses $\Omega_{p,c}(t)$ and the detuning $\Delta_{0}$(in the absence of any error) are applied such that states $|01\rangle$, $|10\rangle$, $|11\rangle$ return to their ground states with phases $\phi_{01}$, $\phi_{10}$, $\phi_{11}$, in which $\phi_{01}=\phi_{10}$ by symmetry. The $C_{Z}$ gate is realized for $\phi_{11}=2\phi_{01}-\pi$ [54]. To realize this gate, we resort to the protocol using double sequences of adiabatic STIRAP pulse which provides strongly suppressed insensitivity to the fluctuations in laser amplitude [55]. The adiabatic pulses are given by
$\Omega_{p}(t)$ $=$ $\Omega_{p}^{\max}(e^{-\frac{(t+t_{1})^{2}}{2\omega^{2}}}-e^{-\frac{(t-t_{1})^{2}}{2\omega^{2}}})$ (3)
$\Omega_{c}(t)$ $=$ $\Omega_{c}^{\max}(e^{-\frac{(t+t_{2})^{2}}{2\omega^{2}}}+e^{-\frac{(t-t_{2})^{2}}{2\omega^{2}}})$ (4)
with $\Omega_{p,c}^{\max}$, $t_{1,2}$, $\omega$ the peak amplitude, the pulse center and the pulse width, respectively. Note that, in addition to the phase change of $\Omega_{p}(t)$ by $\pi$ in the second STIRAP sequence, the sign of intermediate detuning $\Delta_{0}$ is also reversed in the middle of the gate [56], as shown in Fig.1(b).
@@
III.2 Method for numerical calculation
@@ III.3 Pulse optimization towards active robustness
To show a family of Rydberg quantum gates that are actively robust against any detuning error $\epsilon_{\delta}\neq 0$, we start by finding optimal STIRAP pulses $\Omega_{p}(t),\Omega_{c}(t)$ for a two-qubit $C_{Z}$ gate. Here we optimize the gate pulse to be detuning-error robust by employing the prior Genetic algorithm [62] yet containing a cleverly-chosen cost function,
$J_{der}=[1-{{\mathcal{F}}}(\epsilon_{\delta}=0)]^{2}+[{\mathcal{F}}_{\max}(\epsilon_{\delta})-{\mathcal{F}}_{\min}(\epsilon_{\delta})]^{2}$ (14)
Via minimizing this single-objective cost function $J_{der}$ over a range of $\epsilon_{\delta}\in[-\epsilon_{0},\epsilon_{0}]$, we observe that, not only the ideal gate error at $\epsilon_{\delta}=0$ is minimized, but also the sensitivity of the gate fidelity due to the change of $\epsilon_{\delta}$ is significantly reduced, so-called the detuning-error robust(DER) pulse. It is worth emphasizing that such DER pulse has an active robustness to any magnitude of the two-photon detuning error as long as $|\epsilon_{\delta}|<\epsilon_{0}$. However, remember the typical-optimal(TO) pulse only minimizes the ideal gate infidelity in the absence of any error by letting the cost function as
@@
So any deviation $\epsilon_{\delta}\neq 0$ can lead to an exponential increase of infidelity[see Fig.1(c)], in good agreement with Fig.2 of Ref.[35].
@@
To our knowledge the achievable gate fidelity is fundamantally limited by the peak amplitude $\Omega_{p,c}^{\max}$ which also sets a timescale for the gate duration $T_{g}$. Here we fix $\Omega_{p,c}^{\max}/2\pi=150$ MHz, maintaining a high two-photon Rabi frequency $\Omega/2\pi=5.625$ MHz, which restricts $T_{g}$ to be about 2.0 $\mu$s ensured by a minimal pulse area for qubit rotation [63]. Other parameters $\Delta_{0}/2\pi=2.0$ GHz, $B/2\pi=2.0$ GHz are fixed for simplicity. That leaves a few gate parameters $(t_{1},t_{2},\omega)$ to be optimized. To find the DER pulse, we first minimize the cost function $J_{der}$ using a reasonable range for denoting detuning deviation $\epsilon_{0}/2\pi=0.8$ MHz. This realizes an exact $C_{Z}$ gate under optimal pulse parameters
$t_{1}=0.6664\mu s,t_{2}=0.9260\mu s,\omega=0.1666\mu s$ (16)
While the gate parameters for a TO pulse are
$t_{1}=0.4444\mu s,t_{2}=0.9027\mu s,\omega=0.1587\mu s$ (17)
in which the total gate duration is $T_{g}=2(t_{2}+3\omega)=(2.8517,2.3259)$ $\mu$s, a slightly longer due to the use of adiabatic STIRAP pulses. We ignore the intermediate and Rydberg decays by $\gamma_{p,r}=0$ in the optimization and focus on the robustness of the gate pulse against detuning deviation $\epsilon_{\delta}$.
@@
The gate infidelities $1-\mathcal{F}_{der(to)}$ are displayed in Fig.1(c) by the black-solid(DER) and blue-dashed(TO) lines. It is clear that the TO pulse depending on perfect optimization, benefits from a minimal value $\sim 10^{-6}$ only if $\epsilon_{\delta}$ is very small while the infidelity numbers increase exponentially with $\epsilon_{\delta}$ as expected. For a large value of $|\epsilon_{\delta}|/2\pi=1.0$ MHz the infidelity reaches $\sim 0.02$. In contrast, the insensitivity of the DER pulse to the variation of $\epsilon_{\delta}$ is greatly improved although the exact infidelity numbers stay at a relatively high level $10^{-3}\sim 10^{-2}$, yet its robustness is greatly enhanced. The DER pulse for $|\epsilon_{\delta}|/2\pi=0.7$ MHz has an infidelity of 0.0019, smaller than the TO pulse by one order of magnitude. Therefore, the DER optimization has appealing advantages in reducing the sensitivity of gate to the detuning deviation, especially for a larger $\epsilon_{\delta}$ the predicted fidelity limit is still above 0.99. This improvement should be at the expense of a high ideal gate fidelity $\mathcal{F}(\epsilon_{\delta}=0)\sim 99.71\%$(compared to $99.9996\%$ for the TO pulse) because the choice of cost function $J_{der}$ has incorporated the impact of two competing terms (see Appendix B).
Consequently the DER pulse protocol with active robustness will manifest a more dominant merit when the two-photon detuning deviation $\epsilon_{\delta}$ is large. In this case the desirable gate implementation can be made more insensitive to any magnitude of detuning errors, maintaining a stable output performance.
IV DER pulse under different noise sources

### v2

@@ Abstract
Error suppression to the experimental imperfections is a central challenge for useful quantum computing. Recent studies have shown the advantages of using single-modulated pulses based on optimal control which can realize high-fidelity two-qubit gates in neutral-atom arrays. However, typical optimization only minimizes the ideal gate error in the absence of any decay, which allows the gate to be passively influenced by all error sources leading to an exponential increase of the insensitivity when error becomes larger. In the present work, we propose the realization of two-qubit $C_{Z}$ gates with active robustness against two-photon detuning errors. Our method depends on a modified cost function in numerical optimization for shaping gate pulses, which minimizes, not only the ideal gate error but also the fluctuations of gate infidelity over a wide error range. We introduce a family of Rydberg blockade gates with active robustness towards the impacts of versatile noise sources such as Doppler dephasing and ac Stark shifts. The resulting gates with robust pulses can significantly increase the insensitivity to any type of errors acting on the two-photon detuning, benefiting from a relaxed requirement of colder atomic temperatures or more stable lasers for current experimental technology.
@@ I Introduction
The ability to achieve high-fidelity two-qubit quantum gates is a key task for large-scale quantum computing with trapped neutral atoms [1, 2, 3, 4, 5], while noises from versatile experimental imperfections will fundamentally limit the achievable gate fidelity [Nature.626.58(2024)], leading to a huge gap to other state-of-the-art platforms such as superconducting qubits, trapped ions, quantum dots and nuclear magnetic resonance systems [7, 8, 9, 10, 11]. Recent work has demonstrated that the execution of modulated pulses based on optimal control can realize high-fidelity entangling operations [12]. And to date the best reported two-qubit gate fidelity using optimal control and atomic dark state reaches $\mathcal{F}=0.995$ [13]. The optimal control technique enables a smooth modulation of the laser amplitude or phase in time without the requirement for individual addressability of atoms [14, 15, 16, 17, 18, 19]. This outperforms the original $\pi$-gap-$\pi$ method [20, 21, 22] by suppressing severe decoherence between the ground and Rydberg states [22, 23, 24].
@@
The typical optimization focuses on minimizing the ideal gate infidelity in the absence of any noise. e.g. A time-optimal gate contains the merit of minimizing the gate duration together with a smaller time-spent in the Rydberg state which makes the gate intrinsically robust to the decay errors [25]. Whereas, other technical errors will passively affect the gate performance arising the necessity of subsequent error correction ways which is regardless of the system [26, 27, 28, 29, 30, 31]. Recently, performing low-error quantum gates with noise-robust optimization has become a new exciting frontier. To alleviate the insensitivity of gates to type of errors that occur, e.g. constructing a specific cost function that possesses enhanced robustness with respect to some errors, is more attractive [32, 33, 34]. Via utilizing quantum optimal control with GRAPE algorithm [35] a family of Rydberg blockade gates with significantly improved performance is presented that can be simultaneously robust against the amplitude inhomogeneity and the Doppler shifts [36]. Although these works may open unique opportunities for fault-tolerant quantum computing with atomic qubits, preserving an active robustness of systems to type of errors for Rydberg gates is still a big challenge.
@@
Figure 1: Two-qubit $C_{Z}$ gate protocol. (a) Atom-level diagram. Qubits are encoded in two hyperfine clock states $|0\rangle$ and $|1\rangle$ of ${}^{87}\rm{Rb}$ atoms. Atomic excitation to the Rydberg state $|r\rangle$ is carried out by a two-photon transition via a largely-detuned intermediate state $|p\rangle$(detuned by $\Delta_{0}$), driven by 420 nm and 1013 nm lasers. $|d\rangle$ is a leakage state that collects most of the spontaneous decays from $|p\rangle$ and $|r\rangle$ except returning to $|0\rangle$ and $|1\rangle$. The branching radios used for different scattering rates are: $b_{r\to 0}=0.059$, $b_{r\to 1}=0.055$, $b_{r\to d}=0.886$, and $b_{p\to 0}=0.1354$, $b_{p\to 1}=0.2504$, $b_{p\to d}=0.6142$, together with a $3751/\gamma_{r}=$ 375 $\mu$s lifetime for $|r\rangle$ and a $0.1181/\gamma_{p}=$ 0.118 $\mu$s lifetime for $|p\rangle$, chosen from Ref. [37]. The van der Waals interaction leads to a large energy shift $B$ if both atoms are in $|r\rangle$ preventing their simultaneous excitation in the Rydberg blockade limit. $\epsilon_{\Delta}$ and $\epsilon_{\delta}$ represent unknown detuning errors respectively to the intermediate and Rydberg states possibly from versatile experimental imperfections. Bottom of (a) presents typical profiles of the adiabatic pulse sequences $\Omega_{p}(t),\Omega_{c}(t)$ and the non-fluctuated intermediate detuning $\Delta_{0}$. Note that $\Delta_{0}$ is switched between two STIRAP pulses. (b1-b2) The ideal gate infidelity ($\gamma_{p,r}=0$) as a function of the two-photon detuning error $\epsilon_{\delta}$ for different DER and DER-i pulses, where the optimization parameter ranges are $\epsilon_{0}/2\pi=(0.8,0.3)$ and the TO pulse is same.
@@
In this work, we present the realization of two-qubit Rydberg $C_{Z}$ gates with active robustness to all error sources associated with the variation in two-photon detuning [38, 39]. To implement the noise-robust optimization, we modify the cost function in the numerical Genetic algorithm, which minimizes, not only the ideal gate error but also the infidelity fluctuations over a wide error range, enabling the optimal pulse to be actively robust against the variation of detuning errors. In addition, we adopt a realistic model by considering a native two-photon transition using adiabatic double-STIRAP pulse [40] which has shown a better robustness to the variation of experimental parameters such as laser amplitude and pulse shape [41, 42, 43]. To characterize the resulting two-qubit $C_{Z}$ gates, we comparably exploit the results from typical and noise-robust optimizations. The latter can produce experimentally realizable pulses that are more robust to any deviation of two-photon detuning, caused by Doppler dephasing [44] and ac Stark shift [45, 46]. Our approach can be easily generalized to other type of errors paving the avenue for realizing noise-robust quantum gates in future experiments.
@@
Finally we admit that there are two main weakness of this protocol. First, the modified cost function contains two competing terms that means the improvement of robustness should be at the expense of low gate fidelity at small detuning deviations (see Appendix B). For a noise-robust pulse (e.g. the DER pulse), we only achieve the gate with an ideal fidelity of $\mathcal{F}_{der}\approx 0.9971$, depending on a wide range of optimization for enhancing the robustness. While a higher value $\mathcal{F}_{der-i}>0.9999$ is possible by utilizing the fine optimization of the DER-i pulse. Second, the proposed gates will be passively affected by other error sources not associated with the two-photon detuning. As compared to a typical-optimal pulse, our pulses are not explicitly robust to those errors (see Appendix C). Especially, due to the use of an adiabatic pulse sequence that elongates the gate duration, our scheme is more susceptible to the fluctuation in laser phase [47, 48].
II Model description
@@
Consider a level scheme for two-qubit $C_{Z}$ gates, as shown in Fig.1(a). Each atom is modeled as a native four-level system with long-lived hyperfine qubit states $|0\rangle,|1\rangle$, an intermediate state $|p\rangle$ and an uppermost Rydberg state $|r\rangle$ [49]. The coherent excitation from $|1\rangle$ to $|r\rangle$ is mediated by a resonant two-photon transition with laser Rabi frequencies $\Omega_{p},\Omega_{c}$, largely detuned by $\Delta_{0}$ from the intermediate state $|p\rangle$, i.e. $|\Delta_{0}|\gg\Omega_{p},\Omega_{c}$. We assume a strong Rydberg blockade condition where two atoms trapped in two optical tweezers [50] with a distance smaller than the blockade radius, can bring a strong energy shift $B$ to prevent simultaneous excitation of both atoms [51, 52, 53].
@@
for single control or target qubit. Here, $\epsilon_{\Delta}$ and $\epsilon_{\delta}$ treats as the unknown yet constant detuning error over the gate duration, possibly caused by versatile noise sources such as atomic velocity and laser fluctuations. We point out that, the intermediate detuning error $\epsilon_{\Delta}$ is less important because $|\Delta_{0}|\gg\epsilon_{\Delta}$ arising a negligible impact $10^{-10}\sim 10^{-6}$ on fidelity (see Appendix C) [54]. Nevertheless, the laser frequency is resonant with the two-photon transition frequency between $|1\rangle$ and $|r\rangle$, making the role of two-photon detuning error $\epsilon_{\delta}$ more crucial in the scheme.
@@ III.1 Double STIRAP pulse sequences
We proceed by describing the implementation of specific pulse sequences that result in a $\pi$-phase($C_{Z}$) gate. The requirement for a universal controlled-phase gate is that the laser pulses $\Omega_{p,c}(t)$ and the detuning $\Delta_{0}$(in the absence of any error) are applied such that states $|01\rangle$, $|10\rangle$, $|11\rangle$ return to their ground states with phases $\phi_{01}$, $\phi_{10}$, $\phi_{11}$, in which $\phi_{01}=\phi_{10}$ by symmetry. If $\phi_{11}=2\phi_{01}-\pi$ is satisfied it is so-called a $C_{Z}$ gate [55]. To realize this gate, we resort to the protocol by using adiabatic double STIRAP pulse, which in principle provides strongly-suppressed insensitivity to the fluctuations in laser amplitude [56]. The adiabatic pulses are given by
$\Omega_{p}(t)$ $=$ $\Omega_{p}^{\max}(e^{-\frac{(t+t_{1})^{2}}{2\omega_{1}^{2}}}-e^{-\frac{(t-t_{1})^{2}}{2\omega_{1}^{2}}})$ (3)
$\Omega_{c}(t)$ $=$ $\Omega_{c}^{\max}(e^{-\frac{(t+t_{2})^{2}}{2\omega_{2}^{2}}}+e^{-\frac{(t-t_{2})^{2}}{2\omega_{2}^{2}}})$ (4)
with $\Omega_{p,c}^{\max}$, $t_{1,2}$, $\omega_{1,2}$ the peak amplitude, the pulse centers and the pulse widths, respectively. Note that, in addition to the phase change of $\Omega_{p}(t)$ by $\pi$ in the second STIRAP sequence, the sign of intermediate-state detuning $\Delta_{0}$ is also reversed in the middle of the gate [57], as shown in Fig.1a (bottom).
@@
Table 1: Optimized gate parameters for the pulse centers $t_{1},t_{2}$ and pulse widths $\omega_{1},\omega_{2}$. The first pair of DER-i pulses (row 6 and row 7) is associated with Gaussian-type weights while the second pair (row 8 and row 9) is obtained by choosing a set of Uniform-type weights. The total gate duration is solved by $T_{g}=2(t_{2}+3\omega_{2})$. The last column presents the ideal gate fidelity at $\epsilon_{\delta}=0$.
Optimization case Pulse parameters ($\mu$s) Gate performance
[…]
TO $\times$ 0.4277 0.6697 0.1584 0.1584 2.2898 0.9999997
$DER$ 0.8 0.6664 0.9259 0.1666 0.1665 2.8509 0.9970983
$DER$ 0.3 0.2863 0.5332 0.1604 0.1586 2.0180 0.9998366
$DER-i$ 0.8 0.7261 0.9737 0.1619 0.1601 2.9080 0.9995629
$DER-i$ 0.3 0.2874 0.5310 0.1598 0.1588 2.0130 0.9999807
$DER-i$ 0.8 0.7285 0.9788 0.1619 0.1631 2.9363 0.9991733
$DER-i$ 0.3 0.2722 0.5196 0.1600 0.1591 1.9938 0.9998411
III.2 Numerical calculation method
@@ III.3 Pulse optimization towards active robustness
The DER pulse.- To show a family of Rydberg quantum gates that are actively robust against any detuning error, we start by finding optimal STIRAP pulses for a two-qubit $C_{Z}$ gate. Here we optimize the gate pulses to be detuning-error robust with numerical Genetic algorithm as in our prior work [63], yet replaced by using a cleverly-chosen cost function,
$J_{der}=[1-{{\mathcal{F}}}(\epsilon_{\delta}=0)]^{2}+[{\mathcal{F}}_{\max}(\epsilon_{\delta})-{\mathcal{F}}_{\min}(\epsilon_{\delta})]^{2}.$ (14)
Via minimizing this single-objective cost function $J_{der}$ over the range of $\epsilon_{\delta}\in[-\epsilon_{0},\epsilon_{0}]$ ($\epsilon_{0}$ is assumed to be the maximal deviation), we observe that, not only the ideal gate error at $\epsilon_{\delta}=0$ is minimized, but also the insensitivity of gate due to the change of $\epsilon_{\delta}$ is significantly improved, so-called the detuning-error robust (DER) pulse. It is worth emphasizing that, we expect such a DER pulse has an active robustness to any magnitude of the two-photon detuning error $\epsilon_{\delta}$ as long as $|\epsilon_{\delta}|<\epsilon_{0}$ is satisfied. Actually, there exists a balance between the fidelity at zero detuning (or small errors) and the robustness to the change of $\epsilon_{\delta}$ [64]. Based on the choice of $J_{der}$, increasing the robustness should be at the expense of ideal gate fidelity (for more discussions, see Appendix B). In contrast, the typical-optimal (TO) pulse only minimizes the ideal gate infidelity in the absence of any error by letting the cost function simply as
@@
Any deviation $\epsilon_{\delta}\neq 0$ for the TO pulse can lead to a dramatic increase of the infidelity numbers [see Fig.1(b1-b2), the blue-dashed line], in good agreement with Fig.2 of Ref. [36].
@@
To our knowledge the achievable gate fidelity is fundamentally limited by the peak amplitude $\Omega_{p,c}^{\max}$ which also sets a timescale for the gate duration $T_{g}$. Here we fix $\Omega_{p,c}^{\max}/2\pi=150$ MHz, maintaining a high two-photon Rabi frequency $\Omega/2\pi=5.625$ MHz, which restricts $T_{g}$ to be around $2.0$ $\mu$s ensured by a minimal pulse area for qubit rotation [67]. Other parameters $\Delta_{0}/2\pi=2.0$ GHz, $B/2\pi=2.0$ GHz are kept constant throughout the paper. That leaves a few gate parameters $(t_{1},t_{2},\omega_{1},\omega_{2})$ to be optimized. In order to find the DER pulse, we comparably introduce two maximal deviations: $\epsilon_{0}/2\pi=(0.8,0.3)$ MHz, respectively corresponding to $\epsilon_{0}/\Omega=(14.2\%,5.3\%)$ for showing the power of optimization algorithm under different deviation degrees. We minimize the cost function $J_{der}$ for each $\epsilon_{\delta}$ within the range of $\epsilon_{\delta}\in[-\epsilon_{0},\epsilon_{0}]$, which realizes an exact $C_{Z}$ gate with optimal pulse parameters shown in Table 1. Note that we ignore the intermediate and Rydberg decays by $\gamma_{p,r}=0$ in the optimization, and focus on the robustness of the gate pulse against detuning deviations.
@@
The gate infidelities $1-\mathcal{F}_{der(to)}$ as a function of the detuning deviation $\epsilon_{\delta}$, are displayed in Fig.1(b1-b2) by the black-solid (DER) and blue-dashed (TO) lines, corresponding to $\epsilon_{0}/2\pi=0.8$ MHz and 0.3 MHz. It is clear that the TO pulse depending on ideal optimization, benefits from a minimal infidelity $1-\mathcal{F}_{to}<10^{-6}$ only if $\epsilon_{\delta}$ is very small while the infidelity numbers increase exponentially with $\epsilon_{\delta}$. For a large value of $|\epsilon_{\delta}|/2\pi=0.8$ MHz the infidelity even reaches more than $0.02$. On the contrary, we excitingly find that the insensitivity of the robust DER pulse against the variation of $\epsilon_{\delta}$ has been greatly improved. See (b1), it is remarkable that the DER pulse becomes strongly robust against detuning deviations although the exact infidelity numbers stay at a relatively high level $10^{-3}\sim 10^{-2}$. E.g. at $|\epsilon_{\delta}|/2\pi=0.7$ MHz, the DER pulse has an infidelity of 0.00195, smaller than the TO pulse by one order of magnitude. Thus, for a wide optimization range, the DER pulse can reveal appealing advantages in improving the insensitivity of gate to detuning deviations, which may be useful for preserving the predicted gate fidelities above 0.99 for a larger error. But, this improvement should be at the expense of low ideal gate fidelity $\mathcal{F}_{der}(\epsilon_{\delta}=0)\approx 0.9971$ (compared to $\mathcal{F}_{to}\approx 0.9999997$ for the TO pulse) because the cost function $J_{der}$ has incorporated the impact from two competing terms (see Appendix B).
Accounting for the fact that a real deviation $\epsilon_{\delta}$ is not large enough after considering various experimental imperfections, thereby we next use $\epsilon_{0}/2\pi=0.3$ MHz to perform a fine numerical optimization. The results are displayed in Fig.1(b2). Here, in order to achieve the robust DER pulse we find the gate duration $T_{g}$ tends to be shortened by diminishing the wait-time between two STIRAP sequences which leads to a larger gate fidelity $\mathcal{F}_{der}(\epsilon_{\delta}=0)\approx 0.9998$. Moreover, as $\epsilon_{\delta}$ increases, the DER pulse outperforms the TO pulse over a wider range of deviations $\epsilon_{\delta}/2\pi>0.114$ MHz because the time-spent in the Rydberg states has been decreased with a fine-optimization. Consequently, we remark that the DER pulse with active robustness can manifest a more promising merit when the two-photon detuning deviation $\epsilon_{\delta}$ is large, while the TO pulse performs better only if all imperfections are very small. Using the DER pulse enables the gate implementation to be more insensitive against any magnitude of detuning errors, maintaining a stable output performance.
The DER-i pulse.- We identify that the DER pulse has promising advantages for larger two-photon detuning errors. However, maintaining this advantage in realistic experimental conditions requires the full knowledge of noise types. As a Doppler dephasing error on the Rydberg state contributing to detuning errors, typically follows a standard Gaussian distribution [65], we further modify the cost function to be $J_{der-i}$ in the numerical optimization for improving its relevance on the choice of noise type,
The first term in Eq.(16) suggests maximizing the ideal gate fidelity in the absence of any noise, same as $J_{der}$. While the second term means the average gate fidelity ¯$\bar{\mathcal{F}}(\epsilon_{\delta})$ should be simultaneously maximized, where [66]
with variables $\epsilon_{\delta i}$ uniformly extracted from $[-\epsilon_{0},\epsilon_{0}]$. Specifically, it is worth noting that we consider a set of Gaussian-type weights $\omega_{1},\omega_{2},...,\omega_{n}$ with $\sum_{i=1}^{n}\omega_{i}=1$ to define the average fidelity ¯$\bar{\mathcal{F}}(\epsilon_{\delta})$. Within an arbitrary deviation range $[-\epsilon_{0},\epsilon_{0}]$, we solve a large number of discrete fidelity numbers $\mathcal{F}(\epsilon_{\delta i})$ alongside with a Gaussian-weighted random coefficient $\omega_{i}$ which finally leads to a more realistic estimation for the average fidelity ¯$\bar{\mathcal{F}}(\epsilon_{\delta})$. Such setting of $J_{der-i}$ contains a tunable and noise-dependent weight $\omega_{i}$ that can be easily applied to other noise sources.
The ideal infidelity $1-{\mathcal{F}}_{der-i}$ as a function of the detuning error $\epsilon_{\delta}$ with optimized DER-i pulses, is displayed in Fig.1(b1-b2) by the red-dotted lines. We identify that the robust DER-i pulse always outperforms the DER pulse by more than one order of magnitude for small detuning errors, and even satisfies $\mathcal{F}_{der-i}>0.9999$ (ideal fidelity) due to the fine optimization. This is expected because the Gaussian sampling of $\epsilon_{\delta}$ in defining ¯$\bar{\mathcal{F}}(\epsilon_{\delta})$ intrinsically makes the real detuning error felt by atoms smaller, hence the gate fidelity is relatively large when $\epsilon_{\delta}$ is small. The DER-i pulse tends to achieve a larger value of fidelity for small detuning errors, consequently the robustness of which suffers from a small decrease as compared with the DER pulse. But now the DER-i pulse becomes the best choice for reasonable detuning errors by taking account of both a large gate fidelity and the robustness to detuning imperfections. We find it can achieve an infidelity of $1-\mathcal{F}_{der-i}<10^{-4}$ even for $|\epsilon_{\delta}|/2\pi$ up to 0.1 MHz when the optimization range is fine. As for a wide optimization case, the robustness of DER-i pulse can also be improved by restricting the fluctuation of infidelity numbers within $10^{-4}\sim 10^{-2}$ for any $\epsilon_{\delta}$, significantly better than the TO pulse.
IV Robust pulses under different noise sources